OpenStudy (anonymous):

A rectangular pen with a perimeter of 60FT is needed to house three pigs. Determine an equation to model the area of this pen

OpenStudy (anonymous):

x*y=a 2x+2y=60 (in case you needed perimeter equation)

OpenStudy (anonymous):

I was thinking to use A=LxW A= x (60-x) A= 60x - x^2

OpenStudy (anonymous):

x and y would be the length and width, sorry.

OpenStudy (anonymous):

but you would need to use 60-2x, because the perimeter contains 2 widths and 2 lengths.

OpenStudy (anonymous):

So what would be the equation to find the area

OpenStudy (anonymous):

you had it right, you would just need to change x(60-x) to x(60-2x) then simplify

OpenStudy (anonymous):

how would i simply it

OpenStudy (anonymous):

simplify *

OpenStudy (anonymous):

\[2x+2y = 60, x+y =30, x=30-y\] \[A=length*width=y*(30-y) = 30y-^{2}\]